OpenStudy (anonymous):

Find the solution to the equation 64^(x – 4) = 4^2x.

OpenStudy (maheshmeghwal9):

use log.

OpenStudy (anonymous):

I knoww but after that i put it in log form i don't know what to do after.

OpenStudy (anonymous):

Im not very good at Logs

OpenStudy (maheshmeghwal9):

to know about logs go to closed questions @nyan21

OpenStudy (kropot72):

Take logs of both sides: \[(x-4)\ln 64=2x \ln 4\] \[\frac{\ln 64}{\ln 4}=\frac{2x}{x-4}\]

OpenStudy (turingtest):

or write 64 as a power of 4 and use exponent properties\[64^{x-4}=(4^3)^{x-4}=4^{2x}\]now use the rule\[(x^a)^b=x^{ab}\]to simplify what you have and take log base 4 of both sides